The index of refraction of fused quartz is 1.472 for light of wavelength 400 nm and is 1.452 for light of wavelength 760 nm. Find the speeds of light of these wavelengths in fused quartz.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. We know that, μ base 2/ μ base 1 = v base 1/v base 2 So, 1472/1 = 3 * 10^8/v base 400 ⇒ 2.04 * 10^8 m/sec. [because, for air, μ = 1 and v = 3 * 10^8 m/s] Again, 1452/1 = 3 * 10^8/v base 760 ⇒ v base 760 = 2.07 * 10^8 m/sec.